# Thermodynamics and dynamics of coupled complex SYK models

### v3

@@ Abstract
It has been known that the large-$q$ complex SYK model falls under the same universality class as that of van der Waals (mean-field) and saturates the Maldacena-Shenker-Stanford bound, both features shared by various black holes. This makes the SYK model a useful tool in probing the fundamental nature of quantum chaos and holographic duality. This work establishes the robustness of this shared universality class and chaotic properties for SYK-like models by extending to a system of coupled large-$q$ complex SYK models of different orders. We provide a detailed derivation of thermodynamic properties, specifically the critical exponents for an observed phase transition, as well as dynamical properties, in particular the Lyapunov exponent, via the out-of-time correlator calculations. Our analysis reveals that, despite the introduction of an additional scaling parameter through interaction strength ratios, the system undergoes a continuous phase transition at low temperatures, similar to that of the single SYK model. The critical exponents align with the Landau-Ginzburg (mean-field) universality class, shared with van der Waals gases and various AdS black holes. Furthermore, we demonstrate that the coupled SYK system remains maximally chaotic in the large-$q$ limit at low temperatures, adhering to the Maldacena-Shenker-Stanford bound, a feature consistent with the single SYK model. These findings establish robustness and open avenues for broader inquiries into the universality and chaos in complex quantum systems. We provide a detailed outlook for future work by considering the “very” low-temperature regime, where we discuss relations with the Hawking-Page phase transition observed in the holographic dual black holes. We present preliminary calculations and discuss the possible follow-ups that might be be taken to make the connection robust.
@@ I Introduction
There is a caveat, though. Only for the uncharged case does the smaller black hole evaporate, such that we are left with regular (non-chaotic) non-interacting thermal radiation for a particular value of temperature [9]. On the SYK side, however, there exists a whole range of temperatures where we have a chaotic to regular (non-chaotic phase transition) [4] (see section VI.2 below). Although the thermodynamics between the SYK model and thermodynamically holographic dual does match, there is a short-coming of this duality that it does not reproduce the standard Hawking-Page transition [9]. We make this point precise as outlook for future work in sections VI.2 and VI.3 below. We provide preliminary calculations in the outlook and highlight what needs to be done as a future work to make this connection robust, if at all possible. But at the same time, the SYK model has proved instrumental in exploring connections between quantum chaos and black holes [10, 11]. The natural question to ask is how robust are these quantum chaotic and thermodynamic properties of SYK-like models.
@@
In this paper, we study the robustness of the chaotic and thermodynamic properties of a system of coupled large-$q$ complex SYK models and the associated dynamical inequivalence to black holes in the context of Hawking-Page transition. We do this by considering the sum of two non-commuting SYK models of different orders. This has the effect of introducing an additional scale. As expected from relevance, in a renormalization context, the lower order term dominates at low temperatures. Note that typical physical systems will always have at least two competing terms, one associated with its kinetic energy, the other with particle interactions. A strongly dominating kinetic term results in a free, hence non-chaotic, system. As such, it seems reasonable, that such a term would have a large effect on the thermodynamics. Here, interactions are of order $q$, while we add a pseudo-kinetic term of order $q/2$. The term “pseudo-kinetic” is motivated by various findings. For one, such a term coupling a lattice of dots at equilibrium yields the standard intermediate temperature resistivity related to the $q=4$, hence quadratic hopping $q/2=2$, case [12, 13]. This is in agreement with the observation that the large-$q$ expansions converge rather fast, leading to a large quantitative and qualitative overlap even with the $q=4$ case as discussed at length in [14]. This then gives us the advantage of an analytically solvable system via a $1/q$ expansion, as opposed to the $q=4$ case which is only numerically solvable. This is our motivation in calling the $q/2$ term in our model below as “pseudo-kinetic” term because of this convergence and qualitative similarity with the finite $q$ case, such as $q=4$ while retaining the analytical tractability of the system. However we note that since the $q/2$ model is also chaotic for values larger than four, as opposed to the non-chaotic kinetic ($q=2$) SYK model, therefore we expect the term “pseudo-kinetic” to be less appropriate at low temperatures. Another difference that occurs is that the large-$q$ case thermalizes instantly [15] while the $q=4$ case thermalizes at the Planckian rate [16]. The Majorana variant (corresponding to the uncharged case) of our model was considered in [17]. However, as we will see in this work that including a $U(1)$ symmetry yields a much richer phase diagram where we observe a phase transition at low-temperatures.
@@ I.1 Results
We introduce the model and the associated setup in Section II where we present the exact solutions of Green’s functions. Section III discusses the thermodynamics of our model where we evaluate the equation of state as well as the grand potential. We observe a continuous phase transition at low-temperature. We analyze the associated critical point in section IV where we explicitly evaluate the critical exponents and find them to belong to the same universality class as Landau-Ginzburg mean-field exponents, a universality class also shared by van der Waals fluids and various AdS black holes. Then we analyze the dynamics of our model in section V where we study the out-of-time correlator (OTOC) and calculate the Lyapunov exponent. We find that our system is chaotic and saturates the MSS bound at low temperatures. We also discuss the relation to the phase transition at “very” low-temperature (we define therein what “very” low temperatures are) and the associated comparison with the Hawking-Page phase transition. We conclude and provide an outlook in section VI.
@@ II Model
where $\mathcal{T}_{\mathcal{C}}$ is the time ordering operator on the Keldysh contour [19]. This Green’s function may be separated into the greater ($t_{1}>t_{2}$) and lesser ($t_{1}>t_{2}$) Green’s functions, which depict the order of the operators. We use the ansatz for the Green’s functions of large-$q$ complex SYK model [20, 4, 18]
${\cal G}^{\gtrless}(t_{1},t_{2})=({\cal Q}\mp 1/2)e^{g^{\gtrless}(t_{1},t_{2})/q}$ (5)
with the boundary conditions
@@
$\beta\alpha=-2\sqrt{\frac{\left(\beta{\cal J}_{q/2}\right)^{4}}{\pi^{2}v^{2}}+\left(\beta{\cal J}_{q}\right)^{2}}\sin(\pi v/2)$ (17)
@@ III.3 Grand potential
We proceed to calculate the dimensionless contribution $q\beta\Omega$ which is of order ${\cal O}(q^{0})$. We label this as $q\beta{\tilde{\Omega}}$ where we subtract off the logarithmic contribution to the entropy, namely $\tilde{\beta}{\tilde{\Omega}}=q\beta\Omega+q\ln 2$, thereby leaving us with (recall $\tilde{{\cal Q}}\equiv\sqrt{q}{\cal Q}={\cal O}(q^{0})$ in eq. (19))
@@
Next we use the equation of state at low-temperature given in eq. (28) combined with the scaling for $\beta$ and $\mu$ in eqs. (20) and (21), respectively. The resulting expression for $q\beta\mu{\cal Q}$ is of the order ${\cal O}(q^{0})$ and is given by
@@ V Dynamics and chaos
In this section, we will compute the Lyapunov exponent for our coupled system of complex large-$q$ SYK models and show that the model saturates the Maldacena-Shenker-Stanford (MSS) bound at low-temperature [23], thereby being maximally chaotic to order ($\order{1/q}$ ) in the low-temperature limit.
@@
where the indices $j,k$ are the fermionic sites, and $\mathcal{F}=\sum_{n}\mathcal{F}_{n}$ is the unregularized out-of-time-order correlator (OTOC). The first term is a disconnected diagram, and we are interested in the ($\order{\frac{1}{N}}$ ) diagrams. Here, the disorder average is non-vanishing if the number of indices for the two couplings $X_{\mu_{1}\dots\mu_{\kappa q/2}}^{\nu_{1}\dots\nu_{\kappa q/2}}$ (where $\kappa=\{\frac{1}{2},1\}$) are the same. As a result, the average over the product of the $q$ interaction coupling with the $q/2$ interaction is always zero. As an example, take $q=4$, for which we have the disorder average $\expectationvalue{X_{\mu_{1}\mu_{2}}^{\nu_{1}\nu_{2}}\;X_{\mu_{1}}^{\nu_{1}}}=0$.
@@
Figure 3: The diagrams related to the ($\order{\frac{1}{N}}$ ) part of the correlator for our system of coupled complex large-$q$ SYK models. The diagram is drawn for $q=8$, and also consists of a diagram with $\tau_{3}\leftrightarrow\tau_{4}$, which have been omitted here. There are $q-2$ lines connecting the two horizontal rails, out of which half of them ($q/2-1$ lines) run in one direction while the remaining half ($q/2-1$ lines) run in the opposite.
@@
Figure 4: In the large-$N$ limit, the diagrams of order ($\order{\frac{1}{N}}$ ) in the out-of-time-correlator (OTOC) for the coupled SYK system can be constructed with a kernel $K$. The OTOC is then written as $\mathcal{F}=\mathcal{F}_{0}+K\mathcal{F}$, where $\mathcal{F}=\sum\limits_{n=0}^{\infty}\mathcal{F}_{n}$, where $\mathcal{F}_{0}$ is defined in Fig. 3 and becomes suppressed at large time due to exponential growth of the OTOC. This figure shows the process $\mathcal{F}_{1}=K\mathcal{F}_{0}$ for $q=8$, where the term in parentheses is the kernel.
@@ V.2 Chaos regime
[…]
@@ V.3 The Lyapunov exponent
where $\lambda_{L}$ is the growth exponent, known as the Lyapunov exponent. With the ansatz (61), and assuming the function $\mathcal{F}(t_{1},t_{2})$ decays at the boundaries $t_{1},t_{2}\to-\infty$, this reduces to a second order differential equation
@@ V.3.1 Perturbation theory
[…]
@@ VI.1 Conclusion
We now proceed to discuss the “very” low-temerpature regime, where we can make connections with the Hawking-Page phase transition on the holographically dual black holes. We present preliminary calculations and clearly highlight the steps that can be taken as future work to make the connection robust with the first-order Hawking-Page transition, if at all possible.
@@ VI.2 Relation to the phase transitions at “very” low-temperature
(equivalently $\bar{\beta}=\beta q^{-2}$) which is the regime one must consider to carefully take the zero temperature limit. The chemical potential scaling is defined as $\mu\equiv\bar{\mu}q^{2}$. In this regime, much like the standard $q$-body SYK model, the “gaseous” and “liquid” densities diverge [4]. The terminologies are used in accordance with the van der Waals gas as we have already seen that both belong to the same universality class. This is highlighted in Fig. 5. In fact, the liquid phase, the blue line in Fig. 1 (also see Fig. 5), is at $Q_{\ell}={\cal O}(q^{0})$ while the gaseous phase has ${\cal Q}_{g}={\cal O}(q^{-1})$. Now recall that the effective SYK interactions have a Gaussian suppression in charge density. As such, the large charge density of the liquid eventually fully suppresses the SYK interactions, leaving a free non-interacting liquid
@@
One might consider the case that such an additional rescaling of temperature as in eq. (76) might invalidate the leading order in $q$ result given by $g(t)$. By this we mean that writing the Green’s function ${\cal G}(t)\propto e^{g(t)/q+g_{2}(t)/q^{2}+{\cal O}(1/q^{3})}$, the question is whether $g(t)$ and $g_{2}(t)$ can both contribute to the leading order results in the equation of state or the grand potential. At least for the single SYK dot, with $q$-body interactions, this was shown not to be the case. This was shown in appendix A of [4]. In particular, for any polynomial temperature $T={\cal O}(q^{\alpha})$, where $\alpha\leq 0$, only the leading order correction $g(t)$ contributes to the thermodynamics. Naturally, just because this is true in a single large-$q$ SYK model does not guarantee that the same will hold for our coupled system. The study of this correction term would be one on its own, and whether it has any interesting effects on the thermodynamics is left as a future work.
@@
Figure 6: Phase diagram for our coupled SYK system for any finite value of $\gamma$ at low and “very” low temperatures, as discussed in section VI.2. See the text below eq. (81) for the assumption involved in the “very” low-temperature regime that is left as future work.

### v4

@@ Abstract
It has been known that the large-$q$ complex SYK model falls under the same universality class as that of van der Waals (mean-field) and saturates the Maldacena-Shenker-Stanford bound, both features shared by various black holes. This makes the SYK model a useful tool in probing the fundamental nature of quantum chaos and holographic duality. This work establishes the robustness of this shared universality class and chaotic properties for SYK-like models by extending to a system of coupled large-$q$ complex SYK models of different orders. We provide a detailed derivation of thermodynamic properties, specifically the critical exponents for an observed phase transition, as well as dynamical properties, in particular the Lyapunov exponent, via the out-of-time correlator calculations. Our analysis reveals that, despite the introduction of an additional scaling parameter through interaction strength ratios, the system undergoes a continuous phase transition at low temperatures, similar to that of the single SYK model. The critical exponents align with the Landau-Ginzburg (mean-field) universality class, shared with van der Waals gases and various AdS black holes. Furthermore, we demonstrate that the coupled SYK system remains maximally chaotic in the large-$q$ limit at low temperatures, adhering to the Maldacena-Shenker-Stanford bound, a feature consistent with the single SYK model. These findings establish robustness and open avenues for broader inquiries into the universality and chaos in complex quantum systems. We provide a detailed outlook for future work by considering the “very” low-temperature regime, where we discuss relations with the Hawking-Page phase transition observed in the holographic dual black holes. We present preliminary calculations and discuss the possible follow-ups that might be taken to make the connection robust.
@@ I Introduction
There is a caveat, though. Only for the uncharged case does the smaller black hole evaporate, such that we are left with regular (non-chaotic) non-interacting thermal radiation for a particular value of temperature [9]. On the SYK side, however, there exists a whole range of temperatures where we have a chaotic to regular (non-chaotic phase transition) [4] (see section VI.2 below). Although the thermodynamics between the SYK model and thermodynamically holographic dual does match, there is a short-coming of this duality in that it does not reproduce the standard Hawking-Page transition [9]. We make this point precise as outlook for future work in sections VI.2 and VI.3 below. We provide preliminary calculations in the outlook and highlight what needs to be done as a future work to make this connection robust, if at all possible. But at the same time, the SYK model has proved instrumental in exploring connections between quantum chaos and black holes [10, 11]. The natural question to ask is how robust are these quantum chaotic and thermodynamic properties of SYK-like models.
@@
In this paper, we study the robustness of the chaotic and thermodynamic properties of a system of coupled large-$q$ complex SYK models and the associated dynamical inequivalence to black holes in the context of Hawking-Page transition. We do this by considering the sum of two non-commuting SYK models of different orders. This has the effect of introducing an additional scale. As expected from relevance, in a renormalization context, the lower order term dominates at low temperatures. Since an action consists of a kinetic and a potential term, a strongly dominating kinetic term results in a free, hence non-chaotic, system. As such, it seems reasonable, that such a term would have a large effect on the thermodynamics. Here, interactions are of order $q$, while we add a pseudo-kinetic term of order $q/2$. The term “pseudo-kinetic” is motivated by various findings. For one, such a term coupling a lattice of dots at equilibrium yields the standard intermediate temperature resistivity related to the $q=4$, hence quadratic hopping $q/2=2$, case [12, 13]. This is in agreement with the observation that the large-$q$ expansions converge rather fast, leading to a large quantitative and qualitative overlap even with the $q=4$ case as discussed at length in [14]. This then gives us the advantage of an analytically solvable system via a $1/q$ expansion, as opposed to the $q=4$ case which is only numerically solvable. This is our motivation in calling the $q/2$ term in our model below as “pseudo-kinetic” term because of this convergence and qualitative similarity with the finite $q$ case, such as $q=4$ while retaining the analytical tractability of the system. However, we note that since the $q/2$ model is also chaotic for values larger than four, as opposed to the non-chaotic kinetic ($q=2$) SYK model; therefore we expect the term “pseudo-kinetic” to be less appropriate at low temperatures. Another difference that occurs is that the large-$q$ case thermalizes instantly [15] while the $q=4$ case thermalizes at the Planckian rate [16]. The Majorana variant (corresponding to the uncharged case) of our model was considered in [17]. However, as we will see in this work that including a $U(1)$ symmetry yields a much richer phase diagram where we observe a phase transition at low-temperatures.
@@ I.1 Results
We introduce the model and the associated setup in Section II where we present the exact solutions of Green’s functions. Section III discusses the thermodynamics of our model, where we evaluate the equation of state as well as the grand potential. We observe a continuous phase transition at low-temperature. We analyze the associated critical point in section IV where we explicitly evaluate the critical exponents and find them to belong to the same universality class as Landau-Ginzburg mean-field exponents, a universality class also shared by van der Waals fluids and various AdS black holes. Then we analyze the dynamics of our model in section V where we study the out-of-time correlator (OTOC) and calculate the Lyapunov exponent. We find that our system is chaotic and saturates the MSS bound at low temperatures. We also discuss the relation to the phase transition at “very” low-temperature (we define therein what “very” low temperatures are) and the associated comparison with the Hawking-Page phase transition. We conclude and provide an outlook in section VI.
@@ II Model
where $\mathcal{T}_{\mathcal{C}}$ is the time ordering operator on the Keldysh contour [19]. This Green’s function may be separated into the greater ($t_{1}>t_{2}$) and lesser ($t_{2}>t_{1}$) Green’s functions, which depict the order of the operators. We use the ansatz for the Green’s functions of large-$q$ complex SYK model [20, 4, 18]
${\cal G}^{\gtrless}(t_{1},t_{2})=({\cal Q}\mp 1/2)e^{g^{\gtrless}(t_{1},t_{2})/q}.$ (5)
where the system has a conserved total charge per flavor defined by
$\mathcal{Q}\equiv\frac{1}{N}\sum_{\alpha=1}^{N}\left[c_{\alpha}^{\dagger}c_{\alpha}-1/2\right].$ (6)
and $g^{\gtrless}$ satisfies the boundary conditions
@@
$\beta\alpha=-2\sqrt{\left(\frac{\left(\beta{\cal J}_{q/2}\right)^{4}}{\pi^{2}v^{2}}+\left(\beta{\cal J}_{q}\right)^{2}\right)}\sin(\pi v/2)$ (18)
@@ III.1 Finding v in the low-temperature limit
$\pi v=\sqrt{\left(\frac{U^{2}}{q^{2}}+\left(\frac{K^{2}}{\pi v}\right)^{2}\right)}q^{2}\cos(\pi v/2)+\frac{q^{2}K^{2}}{\pi v},$ (25)
@@ III.3 Grand potential
We proceed to calculate the dimensionless contribution $q\beta\Omega$ which is of order ${\cal O}(q^{0})$. We add a logarithmic contribution $q\ln 2$ in order to cancel the logarithmic contribution of the entropy in Eq. (36). We denote this new expression as $\tilde{\beta}{\tilde{\Omega}}=q\beta\Omega+q\ln 2$ where we follow the standard convention that tilded quantities are of order ${\cal O}(q^{0})$, such as $\tilde{\beta}$ and ${\tilde{\Omega}}$ here. We evaluate this expression to get (recall $\tilde{{\cal Q}}\equiv\sqrt{q}{\cal Q}={\cal O}(q^{0})$ in Eq. (20))
@@
Next, we use the equation of state at low-temperature given in Eq. (29) combined with the scaling for $\beta$ and $\mu$ in eqs. (21) and (22), respectively. The resulting expression for $q\beta\mu{\cal Q}$ is of the order ${\cal O}(q^{0})$ and is given by
@@ V Dynamics and chaos
In this section, we will compute the Lyapunov exponent for our coupled system of complex large-$q$ SYK models and show that the model saturates the Maldacena-Shenker-Stanford (MSS) bound at low-temperature [23], thereby being maximally chaotic to order ${\cal O}{1/q}$ in the low-temperature limit.
@@
where the indices $j,k$ are the fermionic sites, and $\mathcal{F}=\sum_{n}\mathcal{F}_{n}$ is the unregularized out-of-time-order correlator (OTOC). The first term is a disconnected diagram, and we are interested in the ${\cal O}{\frac{1}{N}}$ diagrams. Here, the disorder average is non-vanishing if the number of indices for the two couplings $X_{\mu_{1}\dots\mu_{\kappa q/2}}^{\nu_{1}\dots\nu_{\kappa q/2}}$ (where $\kappa=\{\frac{1}{2},1\}$) are the same. As a result, the average over the product of the $q$ interaction coupling with the $q/2$ interaction is always zero. As an example, take $q=4$, for which we have the disorder average $\langle X_{\mu_{1}\mu_{2}}^{\nu_{1}\nu_{2}}\;X_{\mu_{1}}^{\nu_{1}}\rangle=0$.
@@
Figure 3: The diagrams related to the ${\cal O}{\frac{1}{N}}$ part of the correlator for our system of coupled complex large-$q$ SYK models. The diagram is drawn for $q=8$, and also consists of a diagram with $\tau_{3}\leftrightarrow\tau_{4}$, which have been omitted here. There are $q-2$ lines connecting the two horizontal rails, out of which half of them ($q/2-1$ lines) run in one direction while the remaining half ($q/2-1$ lines) run in the opposite.
@@
Figure 4: In the large-$N$ limit, the diagrams of order ${\cal O}{\frac{1}{N}}$ in the out-of-time-correlator (OTOC) for the coupled SYK system can be constructed with a kernel $K$. The OTOC is then written as $\mathcal{F}=\mathcal{F}_{0}+K\mathcal{F}$, where $\mathcal{F}=\sum\limits_{n=0}^{\infty}\mathcal{F}_{n}$, where $\mathcal{F}_{0}$ is defined in Fig. 3 and becomes suppressed at large time due to exponential growth of the OTOC. This figure shows the process $\mathcal{F}_{1}=K\mathcal{F}_{0}$ for $q=8$, where the term in parentheses is the kernel.
@@ V.2 Chaos regime
$\mathcal{F}(t_{1},t_{2},t_{3},t_{4})={\operatorname{Tr}}\left[y\;c_{j}(t_{1})\,y\,c_{k}^{\dagger}(t_{3})\,y\,c_{j}^{\dagger}(t_{2})\,y\,c_{k}(t_{4})\right],$ (54)
@@ V.3 The Lyapunov exponent
Classically, the Lyapunov exponent measures sensitivity to initial conditions. Mathematically, consider two classical trajectories $\vec{x}(t,\vec{x}_{0})$ and $\vec{x}(t,\vec{x}_{0}+\vec{\delta})$, where $\vec{\delta}$ is infinitesimally small (see Fig. 5). The classical Lyapunov exponent $\lambda_{\text{cl}}$ is defined as:
$\left|\frac{\partial x^{i}(t)}{\partial x^{j}(0)}\right|=\left|\{x^{i}(t),p^{j}(0)\}_{\text{Poisson bracket}}\right|\sim e^{\lambda_{\text{cl}}t}$ (62)
where $x^{i}(t)$ is the $i^{\text{th}}$ component of $\vec{x}(t)$ and $p^{j}(0)$ is the $j^{\text{th}}$ component of conjugate momentum $\vec{p}$ corresponding to $\vec{x}(0)$. If $\lambda_{\text{cl}}>0,$ this indicates that two trajectories, which start from infinitesimally close initial conditions, still diverge exponentially over time.
To generalize this to the quantum case, we use the standard prescription of replacing the Poisson bracket with the commutator where we elevate position $\vec{x}$ and momentum $\vec{p}$ to operators $\hat{X}$ and $\hat{P}$ respectively. In other words, we have
|$\left|\{x^{i}(t),p^{j}(0)\}_{\text{Poisson bracket}}\right|^{2}\to$ $\left\langle\left|\left[\hat{X}(t),\hat{P}(0)\right]\right|^{2}\right\rangle$ (63)
$=\text{TOC}-2{\operatorname{Re}}\left[\text{OTOC}\right]$
[…]
$\text{TOC}=||\rho^{1/2}\hat{X}(t)\hat{P}(0)||^{2}_{F}+||\hat{X}(t)\hat{P}(0)\rho^{1/2}||^{2}_{F}$ (64)
where norm|$\ldots$||| | … | | denotes the Frobenius norm defined by $||A||_{F}\equiv\sqrt{\text{Tr}[A^{\dagger}A]}$. OTOC is the out-of-time-ordered correlator given by:
$\text{OTOC}=\text{Tr}\left[\left(\hat{X}(t)\hat{P}(0)\right)^{2}\rho\right]=\left\langle\left(\hat{X}(t)\hat{P}(0)\right)^{2}\right\rangle$ (65)
Note that we used position and conjugate momentum as representative examples for defining OTOC. This definition can be used for any non-commuting quantum observable. This approach is used to define and expand OTOC for large-$N$ theories in Eq. (51), using the regularized method for calculating OTOC introduced in [23], as employed in Eq. (54) (also see Fig. 2 in [28]). Assuming there is an exponential divergence of OTOC as $e^{\lambda_{L}t}$, this generalizes the classical Lyapunov exponent $\lambda_{\text{cl}}$ to the quantum Lyapunov exponent $\lambda_{L}$, which we now evaluate using the integral in Eq. (61).
Figure 5: Classical chaos is defined by exponential sensitivity to initial conditions: classical trajectories when separated by infinitesimally small distance diverge exponentially with time which defines the classical Lyapunov exponent $\lambda_{\text{cl}}$. The Lyapunov exponent is taken as a measure of classical chaos.
@@
where $\lambda_{L}$ is the growth exponent. With the ansatz (66), and assuming the function $\mathcal{F}(t_{1},t_{2})$ decays at the boundaries $t_{1},t_{2}\to-\infty$, this reduces to a second order differential equation
@@ V.3.1 Perturbation theory
For $A=0$, we are left with the Pöschl-Teller potential $V_{0}(x)=-\text{sech}^{2}(x)$ [29] and the ground state is
$\psi_{0}^{(0)}(x)=\text{sech}(x)/\sqrt{2}.$ (73)
@@ VI.1 Conclusion
We now proceed to discuss the “very” low-temperature regime, where we can make connections with the Hawking-Page phase transition on the holographically dual black holes. We present preliminary calculations and clearly highlight the steps that can be taken as future work to make the connection robust with the first-order Hawking-Page transition, if at all possible.
@@ VI.2 Relation to the phase transitions at “very” low-temperature
(equivalently $\bar{\beta}=\beta q^{-2}$) which is the regime one must consider to carefully take the zero temperature limit. The chemical potential scaling is defined as $\mu\equiv\bar{\mu}q^{2}$. In this regime, much like the standard $q$-body SYK model, the “gaseous” and “liquid” densities diverge [4]. The terminologies are used in accordance with the van der Waals gas, as we have already seen that both belong to the same universality class. This is highlighted in Fig. 6. In fact, the liquid phase, the blue line in Fig. 1 (also see Fig. 6), is at $Q_{\ell}={\cal O}(q^{0})$ while the gaseous phase has ${\cal Q}_{g}={\cal O}(q^{-1})$. Now recall that the effective SYK interactions have a Gaussian suppression in charge density. As such, the large charge density of the liquid eventually fully suppresses the SYK interactions, leaving a free non-interacting liquid
@@
One might consider the case that such an additional rescaling of temperature as in Eq. (81) might invalidate the leading order in $q$ result given by $g(t)$. By this, we mean that writing the Green’s function ${\cal G}(t)\propto e^{g(t)/q+g_{2}(t)/q^{2}+{\cal O}(1/q^{3})}$, the question is whether $g(t)$ and $g_{2}(t)$ can both contribute to the leading order results in the equation of state or the grand potential. At least for the single SYK dot, with $q$-body interactions, this was shown not to be the case. This was shown in appendix A of [4]. In particular, for any polynomial temperature $T={\cal O}(q^{\alpha})$, where $\alpha\leq 0$, only the leading order correction $g(t)$ contributes to the thermodynamics. Naturally, just because this is true in a single large-$q$ SYK model does not guarantee that the same will hold for our coupled system. The study of this correction term would be one on its own, and whether it has any interesting effects on the thermodynamics is left for a future work.
@@
Figure 7: Phase diagram for our coupled SYK system for any finite value of $\gamma$ at low and “very” low temperatures, as discussed in section VI.2. See the text below, Eq. (86) for the assumption involved in the “very” low-temperature regime that is left as future work.